The sides of a triangle are 9, 10, and 14. Find the perimeter of the triangle formed

by joining the midpoints of the three sides.



1
Expert's answer
2021-02-04T07:41:21-0500

It is known that the segment of line joining the middle points of two sides of a triangle is parallel to the third side and its lengths is half the third side's.

Thus, the triangle formed by joining the midpoints of the three sides has sides:

"\frac{9}{2}=4.5,\;\frac{10}{2}=5,\;\frac{14}{2}=7."

And its perimeter: "P=4.5+5+7=16.5."